A 1428 kg car traveling at 73.0 km/h hits a second 1170 kg car traveling at 42.0 km/h in the same direction.  If the first car is traveling at 57.5 km/h after the collision, what is the speed of the second car (in km/h) after the collision?
Answer #1

Conservation of momentum

M1u1 + m2u2 = m1v1 + m2v2

1428×73 + 1170×42 = 1428×57.5 + 1170× v2

v2 = 60.92 km/h